OpenStudy (anonymous):

Awarding Medal. What is the value of the expression begin square root symbol a squared plus b squared minus four end square root symbol minus b squared when a = 6 and b = 2?

OpenStudy (anonymous):

A. -2 B. 0 C. 2 D. 4

OpenStudy (the_fizicx99):

Ehhm hopefully I read it correctly, lol :3 \(\ \sf \sqrt{a^2 + b^2 - 4} -b^2 \) If \(\ \sf \color{blue}{B} = 2\) and \(\ \sf \color{red}{A} = 6 \) Then, \(\ \sf \sqrt{(\color{red}{6})^2 + (\color{blue}{2})^2 - 4}\) - \(\ \sf (\color{blue}{2})^2 \implies \sqrt{36 + 4 - 4} - 4 \implies \sqrt{36} - 4 = 6 - 4 = 2 \)

OpenStudy (the_fizicx99):

There was actually an even faster method of doing this, you could have cancel out the square and taken the square root of 4 and kept the negative sign. (6 + 2 - 2) - 4 = 6 - 4 = 2

OpenStudy (anonymous):

Thank You Very Much I'm So Grateful. <3